Madsen Motorss bonds have 24 years remaining to maturity. Interest is paid annually, they have a $1,000 par value, the coupoq 3
Answer #1

Annual coupon=1000*8%=80

Hence current price=Annual coupon*Present value of annuity factor(9%,24)+$1000*Present value of discounting factor(9%,24)

=80*9.70661177+$1000*0.126404941

=$902.93(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=80[1-(1.09)^-24]/0.09

=80*9.70661177

2.Present value of discounting factor=$1000/1.09^24

=$1000*0.126404941